# TableMind: An Autonomous Programmatic Agent for Tool-Augmented Table Reasoning

### v2

@@ 4.4. Reinforcement Fine-tuning
We first present the multi-objective reward design of RFT, followed by a detailed description of RAPO, to enable the agent to autonomously explore and acquire optimal strategies for both tool interaction and complex reasoning.
4.4.1. Reward Design
We design a multi-objective reward function, composed of three key components—structural correctness, final answer accuracy, and strategic tool interaction—to effectively guide the TableMind agent in exploring tool-use strategies and enhancing the efficiency of RLVR(Yue et al., 2025). To ensure that the agent produces syntactically valid and parsable reasoning trajectories, we define a format reward $R_{\mathrm{format}}$ to enforce structural integrity.
@@ 4.4.1. Reward Design
A binary reward is assigned based on whether the output adheres to the required structured format (correct usage of ¡think¿ and ¡answer¿ tags, with a valid JSON object in the final answer). As the primary indicator of task success, the accuracy reward $R_{\mathrm{acc}}$ directly assesses the correctness of the agent’s final answer against the ground truth. The reward is defined as:
@@
For Table Question Answering, correctness is measured by exact match with ground-truth answers, while for Table Fact Verification, it is measured by label accuracy. To guide efficient and effective tool use, we introduce an auxiliary reward $R_{\mathrm{tool}}$ with an implicit curriculum that encourages exploration early and efficiency later.
@@
where $s$ is the number of global training steps, $\mathbb{I}(\cdot)$ is the indicator function that returns $1$ if at least one tool call in the trajectory is successful; $\beta$ is the positive base reward, $N_{\mathrm{turns}}$ is the number of tool turns, and $C$ is a penalty coefficient. The total reward for each reasoning trajectory is the sum of the above components.
@@ 4.4.2. RAPO
We introduce RAPO, a method to improve group-based policy gradient algorithms. It identifies misaligned trajectories—where the model’s confidence ranking disagrees with the true reward—and boosts their learning signal using rank-aware advantage weighting. The RAPO objective modifies the original GRPO(Shao et al., 2024) framework and incorporates certain techniques from DAPO(Yu et al., 2025). First, following recent work(Hu et al., 2025; Liu et al., 2025), we removes the KL divergence, allowing the model to no longer be constrained by the reference model and thereby expanding the search space. Second, it employs a Token-Level Policy Gradient Loss to reduce the impact of answer length on the gradient. Finally, it uses the Clip-Higher strategy, which promotes system diversity and prevents entropy collapse. The objective function is formalized as:
@@
$\mathrm{clip}(r_{i,t}(\theta),1-\epsilon_{\mathrm{low}},1+\epsilon_{\mathrm{high}})A^{\prime}_{i}\big)\bigg],$
Here, $r_{i,t}(\theta)$ is the probability ratio between the current policy $\pi_{\theta}$ and the old policy $\pi_{\theta_{\text{old}}}$ that was used for data collection, calculated as:
@@
(6) $A^{\prime}_{i}=\gamma_{i}\cdot\frac{R_{i}-\mathrm{mean}(\{R_{j}\}_{j=1}^{G})}{\mathrm{std}(\{R_{j}\}_{j=1}^{G})},$
$A^{\prime}_{i}$ is equal to the base group-normalized advantage multiplied by the rank-aware weighting factor $\gamma_{i}$. For misaligned trajectories, $\gamma_{i}>1$; otherwise, $\gamma_{i}=1$.
The central mechanism of RAPO is its method for identifying and amplifying the learning signal from misaligned trajectory pairs. A misalignment occurs when the model assigns a higher confidence to a low-reward trajectory $o_{l}$ than to a high-reward trajectory $o_{w}$ from the same group. This confidence is measured by the log-probability of the sequence, $\log P(o_{i})$, which is the length-normalized log-probability of its constituent tokens. To detect these misalignments, we introduce a pairwise weighting factor $\gamma_{w,l}$, which acts as a diagnostic flag for each winner loser pair. It checks if the model’s confidence ordering for this specific pair is correct. If the winner’s log-probability does not exceed that of the loser, the pair is considered misaligned, and this factor is elevated to apply extra learning pressure. It is formally defined as:
@@
where $\mathbb{I}[\cdot]$ is the indicator function and $\alpha$ is a hyperparameter controlling the intensity of the re-weighting. Finally, the comprehensive weighting factor $\gamma_{i}$ for an individual trajectory is computed by averaging these pairwise diagnostic flags from all comparisons involving that trajectory. This dynamic, rank-aware weighting scheme allows RAPO to efficiently identify and correct instances where the model’s confidence is poorly calibrated, leading to a more robust and effective policy optimization process.
@@ 4.5. Inference and Evaluation
During the inference phase, the model employs an iterative reasoning framework to solve complex problems. This framework operates as a multi-turn cycle, with each cycle comprising the following core steps. Plan: The model first analyzes the user query and the current state to formulate a strategy for the next action. Action: The model generates executable Python code to query or process tabular data. Execution: A code interpreter securely executes the code and returns the output. Reflection: The model receives the execution result as an observation. It then reflects on this outcome to assess progress, verify correctness, and decide whether to generate the final answer or proceed to the next cycle. This plan-action-reflect loop enables the model to perform multi-step reasoning, leverage intermediate results, and self-correct. The process terminates when the model confirms in its reflection step that a solution has been found, or when a predefined maximum number of cycles is reached. Finally, the model synthesizes all information gathered throughout the cycles to generate the final answer. To evaluate model performance, we use the exact match for table question answering tasks and use accuracy for table fact verification tasks.
@@
Method Models Wikitq TabMWP TabFact
Training-free Tab-CoT (2023)(Jin and Lu, 2023) 60.43 92.21 79.05
PoTable (2024)(Mao et al., 2024) 69.56 95.76 88.93
Chain-of-Table (2024)(Wang et al., 2024c) 68.31 96.17 87.78
Tuning-based TableLlama-7B (2023)(Zhang et al., 2023) 53.32 83.88 82.53
TableGPT2-7B (2024)(Su et al., 2024) 62.46 80.43 78.87
Table-R1 (2025)(Yang et al., 2025) 74.86 96.02 87.17
TableMind 76.82 99.27 91.85
Relative Improvement (+2.61) (+3.22) (+3.28)
@@ 5. Experiments
In this section, we outline the experimental setup of TableMind, including datasets, baselines, and evaluation metrics, followed by an analysis of the results to evaluate its effectiveness.
5.1. Experimental Setup
@@
We evaluate TableMind on three widely used public table reasoning benchmarks: WikiTQ(Pasupat and Liang, 2015), TabMWP(Lu et al., 2022), and TabFact (Chen et al., 2019). WikiTQ focuses on open-domain question answering, where models generate concise lists of short answers by combining information from Wikipedia tables with their related textual context. TabMWP targets mathematical reasoning over tables, involving multi-step numerical calculations that require precise intermediate computation and reasoning over quantitative relationships. TabFact addresses fact verification, determining whether a textual statement is supported or refuted by the structured content of a given Wikipedia table. For training, we construct a mixed dataset by randomly sampling 3,500 instances each from the WikiTQ and TabFact training sets, and 1,000 from TabMWP. All evaluations are conducted on the test sets of the respective datasets to ensure consistency and fair comparison with prior work.
@@
We choose six strong LLM-based methods as baselines for evaluating three table reasoning tasks(Zhao et al., 2024; Ye et al., 2023). Tab-CoT(Jin and Lu, 2023) proposes a method that prompts a large language model to generate a table that models a complex reasoning process in a structured, step-by-step manner. PoTable(Mao et al., 2024) employs a plan-then-execute two-stage framework, where an LLM first generates a complete, executable plan, which is then strictly followed by external tools to solve the tabular task. Chain-of-Table(Wang et al., 2024c) introduces a paradigm centered on Evolving Tables, where the table is updated after each reasoning step and serves as the input for the subsequent operation, thus evolving throughout the reasoning chain. TableLlama(Zhang et al., 2023) fine-tunes the Llama 2 model on a diverse table-instruction dataset named TableInstruct, enabling it to handle complex table tasks without requiring a special model architecture. TableGPT2(Su et al., 2024) introduces an innovative semantic table encoder to capture table structure and content, and utilizes large-scale data for pre-training and fine-tuning to achieve deep understanding and processing of tabular data. Table-R1(Yang et al., 2025) is a reasoning model trained specifically for table tasks using reinforcement learning with verifiable rewards, to enable effective inference-time scaling.
@@
For TableMind, we adopt Qwen3-8B as the backbone model. During the supervised fine-tuning stage, we carefully train on 200 synthetic samples with a learning rate of $1\times 10^{-6}$ for one epoch. In the RFT stage, we explicitly implement the RAPO algorithm using the Verl framework, with vLLM consistently employed for generation. In Eq. (3), the hyperparameters are set as $\rho=0.05$, $C=0.01$, $\beta=0.5$, and the group size $G=8$. The batch size is 128, the learning rate is $1\times 10^{-6}$, the policy temperature is 1, the maximum tool-calling turns per prompt is 3, and the maximum response length for each turn is 2048 tokens. Both training stages are extensively conducted on a 4-GPU A800 cluster. To ensure a fair comparison of generation performance, we standardize the number of samples to 1 across all baselines. For other parameters, we faithfully preserve the original configurations proposed by each method, given that the approaches involve distinct target operations.
@@ 5.2. Table Reasoning Performance
To evaluate the efficacy of TableMind, we conduct a comprehensive comparison against competitive baselines across three diverse benchmarks: the open-domain question answering dataset Wikitq, the tabular mathematical reasoning dataset TabMWP, and the fact verification dataset TabFact. The quantitative results, summarized in Table 1, demonstrate that TableMind achieves state-of-the-art performance, consistently surpassing all other models. In particular, it achieves notable improvements on all three benchmarks compared with the strongest respective baselines.
Unlike Tab-CoT, which relies solely on textual reasoning, PoTable and Chain-of-Table leverage external tools. Offloading deterministic tasks like calculations to these reliable executors minimizes factual errors and directly accounts for their superior performance and robustness. While TableLlama and TableGPT2 are trained with SFT to imitate static reasoning paths, Table-R1 uses a more dynamic RFT framework. This approach allows the model to learn by exploring diverse strategies with reliable feedback, fostering a more robust reasoning capability that surpasses the pattern-matching limitations of SFT and leads to its superior performance. Our approach, TableMind, unifies the strengths of both tool-augmented execution and advanced training paradigms. We leverage external tools for high-fidelity, deterministic operations while concurrently training the model with a RLVR framework. This dual strategy ensures both the factual accuracy of individual steps and the cultivation of a robust, generalizable reasoning capability, allowing our model to surpass prior methods that focus on only one of these aspects.
@@ 5.3. Ablation Studies of Key Components
To determine the optimal strategy for the SFT stage, we conducted experiments on various training dataset sizes and epochs. The results are presented in Table 2. They indicate that the optimal performance is achieved with a sample size of 200 at 1 epoch. While performance scales with the dataset size, increasing the training to 2 epochs degrades results, suggesting overfitting on our relatively small dataset. As illustrated by the reward curve in Figure 3 (a), employing SFT as a warm-up phase enhances the efficiency of the subsequent RFT process. Compared to a pure RFT approach, the SFT-initialized model achieves a significantly higher initial reward score of approximately 1.2 at step 0. As shown in Figure  4, the removal of the SFT phase leads to a degradation in model performance. This finding substantiates the necessity of SFT for our training methodology.
@@
5.3.3. Effect of the Multi-objective Reward Design.
We analyze the impact of our multi-objective reward function, with a particular focus on the contribution of our proposed Strategic Interaction Reward $R_{\text{tool}}$. To do so, we conducted an ablation study by removing this auxiliary component, training an agent that relies solely on the primary accuracy $R_{\text{acc}}$ and format rewards $R_{\text{format}}$. As shown in Figure 4, removing the $R_{\text{tool}}$ term leads to a degradation in performance across all benchmarks. This performance drop highlights that simply rewarding the final correct answer is insufficient for learning complex tool-use strategies. Without the guidance of $R_{\text{tool}}$, the agent tended to generate less efficient and often erroneous reasoning chains. Specifically, the absence of the success incentive and efficiency penalty in $R_{\text{tool}}$ resulted in longer, more redundant tool-call sequences that were less likely to contribute to the correct final answer. The implicit curriculum provided by the reward’s decay weight was also lost, leading to slower convergence. Therefore, the Strategic Interaction Reward proves to be a critical component, not just for improving final accuracy, but for effectively shaping the agent’s behavior towards efficient and strategic problem-solving.
@@ 5.3.4. Effect of RAPO
Ablation studies were performed from two key perspectives to thoroughly assess the effectiveness of our proposed RAPO algorithm. First, we compared the training dynamics of RAPO against the baseline GRPO algorithm. The reward score curves in Figure 3 (b) clearly illustrate this comparison. Throughout the training process, the reward curve for RAPO is consistently and stably higher than that of GRPO. This indicates that RAPO not only converges faster but also exhibits a more stable training process with less volatility, demonstrating its advantages in improving the sample efficiency and stability of reinforcement learning. Second, we quantified RAPO’s contribution to the final task performance by removing it from the full model. As shown in Figure 4, removing RAPO led to a degradation in performance on both benchmark datasets. The performance decreased on both the WikiTQ and TabFact datasets, showing a noticeable drop in Exact Match on WikiTQ and in Accuracy on TabFact. These experimental results validate the effectiveness of RAPO from both the training process and final performance standpoints. RAPO not only optimizes training dynamics for a more efficient and stable learning process, but this advantage also directly translates into tangible improvements in the final task-solving performance.

### v3

@@ 4.4. Reinforcement Fine-tuning
Building upon the warm-up provided by SFT, the RFT stage improves generalization by optimizing the agent with carefully designed reward functions and a policy optimization objective.
4.4.1. Multi-perspective Reward Design
We carefully design a multi-perspective reward function, composed of three key components to effectively guide the TableMind agent in exploring tool-use strategies and enhancing the efficiency of RLVR (yue2025does). To ensure that the agent produces syntactically valid and reasoning trajectories, we define a format reward $R_{\mathrm{format}}$ to enforce structural integrity.
@@ 4.4.1. Multi-perspective Reward Design
A binary reward is assigned based on whether the output adheres to the required structured format. As the primary indicator of task success, the accuracy reward $R_{\mathrm{acc}}$ directly assesses the correctness of the agent’s final answer. The reward is defined as:
@@
For table question, correctness is measured by exact match with ground-truth answers, while for table fact verification, it is measured by label accuracy. To guide efficient and effective tool use, we introduce an auxiliary reward $R_{\mathrm{tool}}$ with an implicit curriculum that encourages exploration early and efficiency later.
@@
where $s$ is the number of global training steps, $\mathbb{I}(\cdot)$ is the indicator function that returns $1$ if at least one tool call in the trajectory is successful, $\beta$ is the positive base reward, $N_{\mathrm{turns}}$ is the number of tool turns, and $C$ is a penalty coefficient. The total reward for each reasoning trajectory is the sum of the above components.
@@ 4.4.2. RAPO
We introduce RAPO, a method to improve group-based policy gradient algorithms. It identifies misaligned trajectories and boosts their learning signal using rank-aware advantage weighting. The RAPO objective modifies the original GRPO (shao2024deepseekmath) framework and incorporates some frontier techniques. First, following recent work (hu2025open; liu2025understanding), we removes the KL divergence, allowing the model to no longer be constrained by the reference model and thereby expanding the search space. Second, it employs a token-level policy gradient loss to reduce the impact of answer length on the gradient. Finally, it uses the clip-higher strategy, which promotes system diversity and prevents entropy collapse. The objective function is formalized as:
@@
$\mathrm{clip}(r_{i,t}(\theta),1-\epsilon_{\mathrm{low}},1+\epsilon_{\mathrm{high}})A^{\prime}_{i}\big)\bigg].$
Here, $r_{i,t}(\theta)$ is the probability ratio between the current policy $\pi_{\theta}$ and the old policy $\pi_{\theta_{\text{old}}}$ that was used for data collection:
@@
(6) $A^{\prime}_{i}=\gamma_{i}\cdot\frac{R_{i}-\mathrm{mean}(\{R_{j}\}_{j=1}^{G})}{\mathrm{std}(\{R_{j}\}_{j=1}^{G})}.$
The central mechanism of RAPO is its method for identifying the learning signal from misaligned trajectory pairs. A misalignment occurs when the model assigns a higher confidence to a low-reward trajectory $o_{l}$ than to a high-reward trajectory $o_{w}$ from the same group. This confidence is measured by the log-probability of the sequence, $\log P(o_{i})$, which is the length-normalized log-probability of its tokens. To detect these misalignments, we introduce a pairwise weighting factor $\gamma_{w,l}$, which acts as a diagnostic flag for each winner loser pair. If the winner’s log-probability does not exceed that of the loser, the pair is considered misaligned, and this factor is elevated to apply extra learning pressure. It is defined as:
@@
where $\mathbb{I}[\cdot]$ is the indicator function and $\alpha$ is a hyperparameter controlling the intensity of the re-weighting.
@@ 4.5. Inference and Evaluation
During the inference phase, the model employs an iterative reasoning framework to solve complex problems. This framework operates in a multi-turn loop, where each iteration consists of several core components. The model first analyzes the user query together with the current state to formulate a strategy for the next action. It then generates executable Python code to query or process the tabular data, which is securely executed by a code interpreter to produce an output. The execution result is subsequently provided to the model as an observation, based on which the model reflects to assess progress, verify correctness, and determine whether to generate the final answer or continue to the next iteration.This planning–action–reflection loop enables the model to perform multi-turn reasoning, leverage intermediate results, and self-correct when necessary. The process terminates either when the model determines during the reflection step that a solution has been reached or when a predefined maximum number of iterations is exceeded. Finally, the model synthesizes all information accumulated throughout the iterations to produce the final answer. For evaluation, we adopt exact match as the metric for table question answering tasks and accuracy for table fact verification tasks.
Table 1. Dataset statistics for training and evaluation. In-domain datasets (TabFact, TabMWP, WTQ) are used for both training and evaluation, while out-of-domain datasets (HiTab, FinQA) are used only for evaluation.
Split TabFact TabMWP WTQ HiTab FinQA
Training 3,500 3,500 1,000
Test (RL) 800 800 800
Evaluation 12,779 7,686 4,344 1,349 1,138
@@
Models In-domain Evaluation Out-of-domain Evaluation
Wikitq TabMWP TabFact HiTab FinQA
Open-source Qwen2.5-72B-Instruct 68.56 97.45 87.34 68.02 35.83
Deepseek-R1 74.63 98.03 86.25 70.37 37.42
Training-free Tab-CoT 60.43 92.21 79.05 60.64 27.81
PoTable 69.56 95.76 88.93 61.96 24.93
Chain-of-Table 68.31 96.17 87.78 64.76 29.47
Tuning-based TableLlama-7B 53.32 83.88 82.53 58.73 15.73
TableGPT2-7B 62.46 80.43 78.87 65.42 17.69
Table-R1 74.86 96.02 87.17 69.24 41.27
TableMind 76.82 99.27 91.85 71.95 42.02
Relative Improvement (+2.61) (+1.26) (+3.28) (+2.24) (+1.82)
@@ 5. Experiments
In this section, we describe the experimental setup for TableMind, including datasets, baselines, and evaluation metrics, and then analyze the results to assess its effectiveness across benchmarks.
5.1. Experimental Settings
@@
Table 1 presents detailed dataset statistics. We employ three in-domain datasets for training and evaluation: WikiTQ, HiTab, and FinQA, constructing high-quality slow-thinking demonstrations using the answer-aware generation approach described in Section 3.2. We evaluate TableMind on comprehensive table reasoning benchmarks, categorized into in-domain and out-of-domain evaluations to fully assess the model’s performance and generalization capabilities. For in-domain evaluation, we utilize three widely used public datasets: WikiTQ (pasupat2015compositional), TabMWP (lu2022dynamic), and TabFact (chen2019tabfact). WikiTQ focuses on open-domain question answering by generating short answers from Wikipedia tables and their associated text.TabMWP targets multi-step mathematical reasoning over tables, requiring accurate numerical computation.TabFact addresses fact verification by judging whether a statement is supported by a given Wikipedia table. To further evaluate the robustness and transferability of TableMind, we extend our experiments to two out-of-domain datasets: HiTab (cheng2021hitab) and FinQA (chen2021finqa). HiTab introduces hierarchical tables that require understanding complex header dependencies for statistical question answering.FinQA targets expert-level financial reasoning, requiring multi-step numerical calculations over both tables and text. All evaluations are conducted on the test sets of the respective datasets to ensure consistency and fair comparison with prior work.
@@
We choose six strong LLM-based methods as baselines for evaluating three table reasoning tasks (zhao2024tapera; ye2023large). Tab-CoT (jin2023tab) proposes a method that prompts a LLM to generate a table that models a complex reasoning process in a structured, step-by-step manner. PoTable (mao2024potable) employs a plan-then-execute two-stage framework, which is strictly followed by external tools to solve the tabular task. Chain-of-Table (wang2024chain) introduces a paradigm centered on evolving tables, where the table is updated after each reasoning step and serves as the input for the subsequent operation, thus evolving throughout the reasoning chain. TableLlama (zhang2023tablellama) fine-tunes the Llama 2 model on a diverse table-instruction dataset named TableInstruct, enabling it to handle complex table tasks without requiring a special model architecture. TableGPT2 (su2024tablegpt2) introduces a semantic table encoder to capture table structure and content, and leverages pre-training and fine-tuning for deep tabular understanding. Table-R1 (yang2025table) is a reasoning model trained specifically for table tasks using reinforcement learning with verifiable rewards, to enable effective inference-time scaling.
@@
For TableMind, we adopt Qwen3-8B as the backbone model. During the supervised fine-tuning stage, we carefully train on 200 synthetic samples with a learning rate of $1\times 10^{-6}$ for one epoch. In the RFT stage, we explicitly implement the RAPO algorithm using the agent-r1 framework, with vLLM consistently employed for generation [^2] https://github.com/0russwest0/Agent-R1. In Eq. (3), the hyperparameters are set as $\rho=0.05$, $C=0.01$, $\beta=0.5$, and the group size $G=8$. The batch size is 128, the learning rate is $1\times 10^{-6}$, the policy temperature is 1, the maximum tool-calling turns per prompt is 3, and the maximum response length for each turn is 2048 tokens. Both training stages are extensively conducted on a 4-GPU A800 cluster. To ensure a strictly fair comparison of generation performance, we standardize the number of samples to 1 across all baselines.
@@ 5.2. Table Reasoning Performance
To evaluate the effectiveness of TableMind, we conduct a comprehensive comparison against competitive baselines across five diverse benchmarks, categorized into in-domain tasks (Wikitq, TabMWP, TabFact) and out-of-domain tasks (HiTab, FinQA). The quantitative results, summarized in Table 2, demonstrate that TableMind achieves the best performance across all five datasets. Notably, it consistently surpasses strong open-source general LLMs and specialized table models across both in-domain and out-of-domain benchmarks. Furthermore, TableMind exhibits strong generalization capabilities on the unseen HiTab and FinQA datasets, highlighting its adaptability to complex structural reasoning. Unlike Tab-CoT, which relies solely on textual reasoning, PoTable and Chain-of-Table leverage external tools to minimize calculation errors, accounting for their improvements over standard CoT. Among tuning-based methods, while TableLlama and TableGPT2 rely on SFT to imitate static reasoning paths, Table-R1 employs a dynamic RFT framework, allowing it to surpass SFT-based limitations. It is worth noting that the strong general reasoning model Deepseek-R1 also shows competitive performance, securing the runner-up position on TabMWP and HiTab. However, our approach, TableMind, unifies the strengths of tool-augmented execution and the RLVR training paradigm. This dual strategy allows TableMind to surpass Deepseek-R1 and Table-R1 on both the complex hierarchical tables of HiTab and the financial reasoning tasks of FinQA.
@@ 5.3. Ablation Studies of Key Components
To determine the optimal strategy for the SFT stage, we conducted experiments on various training dataset sizes and epochs. The results are presented in Table 3. They indicate that the optimal performance is achieved with a sample size of 200 at one epoch. While performance scales with the dataset size, increasing the training to two epochs degrades results, suggesting overfitting on our relatively small dataset. As illustrated by the reward curve in Figure 3 (a), employing SFT as a warm-up phase enhances the efficiency of the subsequent RFT process. Compared to a pure RFT approach, the SFT-initialized model achieves a higher initial reward score during the early stage of training. As shown in Figure  4, the removal of the SFT phase leads to a degradation in model performance. This finding substantiates the necessity of SFT for our training methodology.
@@
5.3.3. Effect of the Multi-perspective Reward Design.
We analyze the impact of our multi-objective reward function, with a particular focus on the contribution of our proposed Strategic Interaction Reward $R_{\text{tool}}$. To do so, we conducted an ablation study by removing this auxiliary component, training an agent that relies solely on the primary accuracy $R_{\text{acc}}$ and format rewards $R_{\text{format}}$. As shown in Figure 4, removing the $R_{\text{tool}}$ term leads to a degradation in performance across all benchmarks. Without the guidance of $R_{\text{tool}}$, the agent tended to generate less efficient and often erroneous reasoning chains. Specifically, the absence of the success incentive and efficiency penalty in $R_{\text{tool}}$ resulted in longer, more redundant tool-call sequences that were less likely to contribute to the correct final answer. The implicit curriculum provided by the reward’s decay weight was also lost, leading to slower convergence.
@@ 5.3.4. Effect of RAPO
Ablation studies were performed from two key perspectives to thoroughly assess the effectiveness of our proposed RAPO algorithm. First, we compared the training dynamics of RAPO against the baseline GRPO algorithm. The reward score curves in Figure 3 (b) illustrate this comparison. Throughout the training process, the reward curve for RAPO is higher than that of GRPO. This indicates that RAPO not only converges faster but also exhibits a more stable training process with less volatility, demonstrating its advantages in improving the sample efficiency of reinforcement learning. Second, we quantified RAPO’s contribution to the final task performance by removing it from the full model. As shown in Figure 4, removing RAPO led to a degradation in performance on both benchmark datasets. The performance decreased on both the WikiTQ and TabFact datasets, showing a noticeable drop in exact match on WikiTQ and in accuracy on TabFact.